A spelunker is surveying a cave. She follows a passage 180 m straight west, then 210 m in a direction 45° east of south, and then 280 m at 30° east of north. After a fourth displacement, she finds herself back where she started. (a) Use a scale drawing to determine the magnitude and direction of the fourth displacement. (b) Use component method to determine the magnitude and direction of the fourth displacement.
1
Expert's answer
2020-08-31T12:51:39-0400

a) See the figure:



Approximately, the magnitude is 150 m, the angle South of West is 45°.

b) Compute x- and y-components of each vector representing displacement:

"1: x=-180,\\
\space \space \space \space \space y=0.\\
2: x=210\text{ sin}45°=148,\\
\space \space \space \space \space y=-210\text{ cos}45°=-148.\\
3: x=280\text{ sin}30°=140,\\
\space \space \space \space \space y=280\text{ cos}30°=242.\\
4: u=?\\
\space \space \space \space \space v=?"

sine she returned to the initial position, the displacement is 0, or x- and y-components of this vector are 0. These components, on the other hand, are the sum of components of vectors 1, 2, 3, and 4. This will help us find u and v:


"Ox:0=-180+148+140+u,\\u=-108,\\
Oy:0=0-148+242+v,\\
v=-94."

The magnitude of vector 4?


"m=\sqrt{v^2+u^2}=143\text{ m}."

The angle (South of West) is


"\theta=\text{atan}\frac{94}{108}=41°."